OpenStudy (anonymous):

Serena earns $60 for 8 hours of work. At this rate how many hours will she have to work to earn $150?

OpenStudy (anonymous):

multiply 60 by 8 then divide by 150?

OpenStudy (anonymous):

20

OpenStudy (anonymous):

that 3.5

OpenStudy (anonymous):

3.2 actully

OpenStudy (anonymous):

60*8=480/150=3.2?

OpenStudy (anonymous):

thats not one of my anwser choices

OpenStudy (anonymous):

post answers